OpenStudy (anonymous):

How will adding the value 90 affect the mean and median of the data set 1, 5, 5, 6, 9, 10? A. The mean increases more than the median increases. B. The mean increases and the median stays the same. C. The median increases and the mean stays the same. D. The mean and the median increase by the same amount.

OpenStudy (anonymous):

Well, your current set is "1, 5, 5, 6, 9, 10" so the mean, or the average, is "6" (sum of all the numbers, then divided by the total amount of numbers) your median is 5 and 6 (1, 5, [5, 6,] 9, 10) so the median would be the average of those two numbers (5+6=11 then divided by 2 is 5.5). when you add 90 to the set, things get wierd! new set is "1, 5, 5, 6, 9, 10, 90" so the new average, or mean, is 18 (1+5+5+6+9+10+90 = 126, 126/2 = 18) and the new median is "6" (1, 5, 5, [6], 9, 10, 90). So, as you can see, your original mean was 6 and your original median was 5.5 your new mean is 18, and your new median is 6. Hope this helps

OpenStudy (anonymous):

So the answer is A? @Alexandermag89

OpenStudy (anonymous):

@Alexandermag89

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

Sorry @VivaLaChloe1, I didn't see your second post until now. A is correct.